OpenStudy (anonymous):

what is the next term in this sequence? 2,6,12,20?

OpenStudy (anonymous):

30?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

2+4=6 6+6=12 12+8=20 20+10=30 ^

OpenStudy (anonymous):

30; the first time u add 4, then 6, then 8, now you add 10, getting 30.